OpenStudy (anonymous):

Use the point slope formula to write an equation of the line given the following. The line passes thru the point (4,-1) and is parallel to the line y=3x+1?

OpenStudy (anonymous):

point slope formula: \[(y+b) =m(x+a)\] where "a" and "b" is your (4,-1) aka (a,b), remember to take the negatives into mind. your slope is the "m" which is the "3" from your original equation which will be the same in your new equation since they are "parallel" your new equation should look similar to this but is not the answer! \[y = 3x + 6\]

OpenStudy (anonymous):

Alright what I have is y-(-1)=3(x-4) y+2=3x-12 and ended with y=3x-13

OpenStudy (anonymous):

Your last constant "13" is a bit incorrect. Remember to subtract the right amount. Otherwise, good job. Sorry for the mistaken sign of "+" and "-" in the first equation. Cheers.

OpenStudy (anonymous):

but wouldn't y-(-1) turn it to y+1? o.o

OpenStudy (anonymous):

you put y+2 as your second output after the first. But you are right y-(-1) is y+1. Lol. The confusion was in your 2nd reply with the "y+2". Sorry, I'm going a bit fast and seem to have not caught the other mistake before realizing the final answer is correct. =)

OpenStudy (anonymous):

Oh good lord I didn't see that I had written 2. So, it ends up being y=3x-13? Is that all I have to do?

OpenStudy (anonymous):

Yes. That's all you have to do. =)

OpenStudy (anonymous):

Gracias